OpenStudy (anonymous):

Which is a better investment: 8 % compounded quarterly or 8.3 % compounded annually?

OpenStudy (anonymous):

8% p.a

OpenStudy (anonymous):

can u explain why

OpenStudy (anonymous):

I am asking 8 pa or per quarter

OpenStudy (anonymous):

8.3% is a tiny bit higher. If you take 8% compounded quarterly, you'll get 8.24% at the end of the year. You can find this by taking (1 + .08/4) ^ 4 = 1.0824, or 8.24%.

OpenStudy (anonymous):

correct if it is 8 pa

OpenStudy (anonymous):

so it is 8% ?

OpenStudy (anonymous):

Compound Interest P is the principal (the money you start with, your first deposit) r is the annual rate of interest as a decimal (5% means r = 0.05) n is the number of years you leave it on deposit A is how much money you've accumulated after n years, including interest. If the interest is compounded once a year: A = P(1 + r)n If the interest is compounded q times a year: A = P(1 + r/q)nq Examples A = 1000(1+0.083)^1 = 1083 (compounded annually) A = 1000(1+0.08/4)^4 = 1082 (compounded quarterly) Better option - 8.3% Yearly